OpenStudy (anonymous):

using L'Hospital's rule, solve lim x->0 cot2x*sin6x and lim x->infinity x-lnx. I know I need common denominators....can't figure out how to get them.

OpenStudy (anonymous):

\[=\lim [\cos(2x)*\sin6x]/\sin(2x)= \[=\lim(-2\sin2x*\sin6x+6\cos2x*\cos6x)/2\cos2x=6/2=3\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos(2x)}{\sin(2x)}*\sin(6x)=\lim_{x \rightarrow0}\frac{\cos(2x)*\sin(6x)}{\sin(2x)}\] we have 0/0 so we can apply lhospital \[=\lim_{x \rightarrow 0}\frac{-2\sin(2x)*\sin(6x)+6\cos(6x)*\cos(2x)}{2\cos(2x)}\]

myininaya (myininaya):

\[=\frac{-2(0)*0+6*1*1}{2*1}=\frac{6}{2}=3\]

OpenStudy (anonymous):

dude u rock. i thought i couldn't do that for some reason, so i'm glad I double checked. good to know i can trust my own instincts about this stuff tho.